OpenStudy (anonymous):

Can you simplify or not simplify this expression (s^a)^m? If you can or can't can somebody please explain why? I am so confused on how to simplify this expression. GIVING MEDAL IF SOMEBODY CAN TELL ME THE CORRECT ANSWER! :D

OpenStudy (zarkon):

here is an example \[(4^2)^3=4^24^24^2=4^{2+2+2}=4^{2\times 3}\]